OpenStudy (anonymous):

simplifty 3/8 /6 * 1/2+1/2 please show work!

OpenStudy (mathteacher1729):

Do you mean \[\frac{\frac{3}{8}}{6}\cdot \frac{1}{2}+\frac{1}{2}\]

OpenStudy (anonymous):

3/8 (/6)*(1/2 +1/2)

OpenStudy (anonymous):

=2.5